OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

not likely

OpenStudy (anonymous):

the expression on the left is negative the number on the right is positive no way to solve such a thing

OpenStudy (anonymous):

i.e. "no solution" clear right? \(10\) is positive but \(-5|x+1|\) is not positive no matter what \(x\) is

OpenStudy (armyrangers):

I would like to point out the absolute value sine @satellite73 so that would make it a positive

OpenStudy (anonymous):

thanks so much, do you think you could help me with another

OpenStudy (armyrangers):

Yeah

OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38

OpenStudy (armyrangers):

Ok so what you are going to do is times two by X and six then take the absolute value then add14to the 12then - the answer by 38 then dived 2X by your other answer then your answer

OpenStudy (anonymous):

2lx - 6l +14 - 14 = 38 - 14 2lx - 6l = 24 2lx - 6l/2 = 24/2 lx - 6l = 12 there will be 2 equations x-6 = 12 x-6 =-12 x = 18 x = -6